OpenStudy (anonymous):

Solve x – 2y = 6 for y. Answer y = x – 3 y = one-halfx – 3 y = negative one-halfx + 3 y = –x + 3

OpenStudy (accessdenied):

isolate the term with y on one side by adding and subtracting the other terms to the opposite side then divide by the coefficient on y.

OpenStudy (anonymous):

y = one-halfx – 3 x – 2y = 6 for y x -6 = 2y ->y = x/2 - 3